OpenStudy (anonymous):

complete the square: x^2+10x+35

OpenStudy (earthcitizen):

(x+5)^2+10

OpenStudy (anonymous):

Please ignore the 35 and use x^ + 10x to get x^ + 10x + 25

OpenStudy (anonymous):

x^2+10x+25+10

OpenStudy (anonymous):

how did you get that @meverett04?

OpenStudy (earthcitizen):

x = -5-i sqrt(10), x = -5+i sqrt(10)

OpenStudy (anonymous):

WTH?

OpenStudy (earthcitizen):

\[x = -5+i \sqrt(10) ,x = -5-i \sqrt(10) \]

OpenStudy (anonymous):

WTH is what im saying right now

OpenStudy (anonymous):

wait, okay so thats the answer but, how did you get that??

OpenStudy (earthcitizen):

it has no real roots, only imaginary roots

OpenStudy (mertsj):